Could you walk me thru this plz?
A circle of diameter 2 in. has its center at the center of a square of side 2 in. Find the area of the region that is inside the square and outside the circle.

Answer by Alan3354(69443)

It's the area of the square minus the area of the circle.
Area of the square = 2*2 = 4 sq units
Area of the circle = pi*r^2 = pi sq units
= 4 - pi
=~ 0.8584 sq units